# SubTrack++ : Gradient Subspace Tracking for Scalable LLM Training

### v2

@@ 4 Experiments and Results
Robust Subspace Tracking. Relying on SVD for subspace updates makes methods sensitive to noise and abrupt changes (He et al., 2025). Figure 5 compares Grassmannian subspace tracking with GaLore’s SVD on the Ackley function, highlighting how SVD causes erratic jumps, while our subspace tracking ensures robust optimization. GaLore struggles to reach the global minimum within 100 steps at scale factor 1, and although increasing the scale factor to 3 improves performance, it amplifies jumps that hinder convergence in non-convex settings, revealing sensitivity to hyperparameters, noise, and abrupt changes.
5 Related Works
Parameter-Efficient Training. Several works aim to improve the efficiency of training LLMs, addressing a growing demand as their popularity rapidly increases. Popular LoRA (Hu et al., 2021) significantly reduces memory requirements for fine-tuning LLMs by leveraging two low-rank trainable low-rank matrices. Dettmers et al. (2024) employ quantization techniques and paged optimizers to further reduce memory usage. Additionally, Yaras et al. (2024) introduce Deep LoRA to address overfitting issues and reducing the need for precise tuning of the rank parameter. Several other works have also extended LoRA to enhance the efficiency of training and fine-tuning LLMs (Lialin et al., 2023; Renduchintala et al., 2024; Xia et al., 2024; Pan et al., 2024). Miles et al. (2024) propose compressing intermediate activations and reconstructing them during backpropagation to enhance memory efficiency. Additionally, Hao et al. (2024) demonstrate that full-parameter fine-tuning is feasible by random projections on the gradient matrix, showing that LoRA essentially performs a down-projection of the gradient. BAdam (Luo et al., 2024) leverages the block coordinate descent framework to reduce memory consumption while maintaining capabilities comparable to Adam.
@@ 5 Related Works
Gradient Low-Rank Projection. Several approaches aim to reduce optimizer states, as optimizers like Adam (Kingma and Ba, 2017) account for a significant portion of memory footprint (Li et al., 2023; Anil et al., 2019; Lv et al., 2024; Dettmers et al., 2022). MicroAdam (Modoranu et al., 2024) tackles this by compressing the gradient space and utilizing the compression error through feedback loops. Adam-mini (Zhang et al., 2024) partitions model into blocks, assigning a single learning rate to each block to preserve performance while saving memory. Gur-Ari et al. (2018); Schneider et al. (2024); Yaras et al. (2023) suggest that a substantial portion of gradients lies within a largely consistent subspace. GaLore (Zhao et al., 2024) leverages this fact to reduce the optimizer’s memory by projecting gradients into a low-rank subspace and then projecting them back for full parameter tuning. This approach has been integrated with other methods regarding efficient LLM training (Li et al., 2024; Rajabi and Rambhatla, 2024b). However, not all layers’ gradients evolve within a stable low-rank subspace. Jaiswal et al. (2024); Rajabi and Rambhatla (2024d) identify layers where gradients evolve within a low-dimensional subspace and fine-tune only those layers, freezing the others to avoid inefficient low-rank updates. Grass (Muhamed et al., 2024) reduces memory usage by applying sparse projection matrices to the gradient. Ramesh et al. (2024) dynamically select and update a subset of parameters, for a fast and memory-efficient training. FIRA (Chen et al., 2025) utilize a norm-based scaling method along with GaLore to maintain performance comparable to full-rank training. GoLore (He et al., 2025) addresses GaLore’s convergence issues and employ random projection in latter steps as a solutions. LDAdam (Robert et al., 2025) performs optimization within lower-dimensional subspaces, incorporating a projection-aware optimization update rule and a generalized error feedback mechanism. APOLLO (Zhu et al., 2025) approximates channel-wise learning rate scaling based on random projection. Also, Liang et al. (2024) introduce a dynamically evolving projection matrix updated via online PCA, enhancing the model’s ability to navigate the parameter space efficiently without relying on expensive SVD, a similar approach is also proposed in Rajabi and Rambhatla (2024a).
@@
Grassmannian and Subspace Tracking. A common approach in working with high-dimensional data is to project the data into a lower-dimensional space, and many studies focus on tracking these subspaces as they evolve. Balzano et al. (2011); Rajabi and Rambhatla (2024c) introduce an incremental method for updating subspaces on the Grassmannian when the data is partially observed. Zhang and Balzano (2016) and Kasai (2017) propose methods to handle noise effect in tracking these subspaces. Furthermore, Blocker et al. (2023) present a method for evolving geodesic-based data in the Grassmannian for updating the subspace effectively.
@@ Appendix A Convergence of SubTrack++
By applying $vec$ to the gradient form given in the theorem, we obtain:
@@ Appendix C Projection-Aware Optimizer
Suppose the subspace changes at step $t$, transitioning from the subspace spanned by the orthonormal matrix $S_{t-1}$ to that spanned by $S_{t}$. Since both matrices are orthonormal—preserved by the Grassmannian update rule in (5)—the matrix $S_{t}^{\top}S_{t-1}$ represents the change of basis between the two subspaces. In other words, if $\mathbb{E}_{t-1}=(e_{t-1}^{1},\dots,e_{t-1}^{r})$ and $\mathbb{E}_{t}=(e_{t}^{1},\dots,e_{t}^{r})$ are orthonormal bases for the subspaces at steps $t-1$ and $t$, respectively, then the $i$th column of matrix $X$ transforms under the change of basis as $X^{i}_{t}=\sum_{j=1}^{r}\langle e^{i}_{t},e^{j}_{t-1}\rangle X^{j}_{t-1}$, where $X^{j}_{t-1}$ is the $j$th column of $X$ based on the basis of the time step $t-1$.
@@ Appendix D Time Complexity Analysis
Table 3 presents the time complexity breakdown for the subspace update step in the SubTrack++ algorithm assuming a $m\times n$ gradient matrix and rank $r$ projection matrix, where $r\ll m\leq n$. As outlined in Algorithm 1, the subspace update step begins by solving the least squares problem (2) to estimate the optimal update for $S_{t}$, the $m\times r$ orthonormal matrix. This operation has a time complexity of $O(mr^{2})$. Computing the residual and the partial derivative with respect to $S_{t}$ requires $O(mrn)$ and $O(mnr)$ time respectively. This is because the solution to the least squares problem, $A$, has shape $r\times n$ which is multiplied by $S_{t}$ in the residual $R=G_{t}-S_{t}A$, resulting in time complexity $O(mrn)$. The following operation for the partial derivative is $-2RA^{T}$, where the matrix multiplication has $O(mnr)$ complexity. The tangent vector computation (4) which involves an identity transformation and matrix multiplication has time complexity of $O(m^{2}r)$.
@@
Rank-$1$ approximation of $\Delta F$ $O(mr^{2})$
@@
Appendix E Fine-Tuning Experiments
As described, we examined SubTrack++ on fine-tuning RoBERTa-base and RoBERTa-large models to evaluate them on GLUE and SuperGLUE benchmarks. The results for GLUE task is summarized in Table 4, and SuperGLUE in 5.
Table 4: Evaluating the performance of SubTrack++ and other baselines when fine-tuning RoBERTa-Base on GLUE tasks for $r=8$. The performance is measured via Accuracy for SST-2 and RTE tasks, F1 for MRPC, Pearson Correlation for STS-B, and Matthews Correlation for COLA. The best results are marked in bold, with the second-best performance underlined.
COLA STS-B MRPC RTE SST-2
Full-Rank 62.57 91.03 91.32 77.98 94.27
BAdam [Luo et al., 2024] 54.44 89.01 91.35 68.59 94.15
GaLore [Zhao et al., 2024] 58.54 90.61 91.30 74.37 94.50
LDAdam [Robert et al., 2025] 58.81 90.90 92.22 76.53 94.27
SubTrack++ (Ours) 58.29 90.90 91.84 76.53 90.83
Table 5: Evaluating the performance of SubTrack++ and other baselines when fine-tuning RoBERTa-Large on SuperGLUE tasks with $r=8$. The performance is measured via Accuracy for COPA, WIC, WSC, BoolQ, and AXg tasks, and F1 for CB. The best results are marked in bold, with the second-best performance underlined.
BoolQ CB COPA WIC WSC AXg
Full-Rank 85.96 90.33 76.00 71.79 63.46 96.30
GaLore [Zhao et al., 2024] 85.44 88.85 80.00 71.47 63.46 100.00
BAdam [Luo et al., 2024] 82.51 53.28 59.00 70.38 60.58 51.85
LDAdam Robert et al. [2025] 85.75 56.16 80.00 71.00 64.42 70.37
SubTrack++ (Ours) 84.68 89.66 88.00 70.06 63.46 100.00
Table 6: Hyperparameters of fine-tuning RoBERTa-Base on GLUE tasks.
SST-2 MRPC CoLA RTE STS-B
Shared Parameters Batch Size 16 16 32 16 16
# Epochs 30
Max Seq. Len. 512
Low-Rank Optimizer Learning Rate 2E-05 2E-05 1E-05 2E-05 3E-05
Methods Parameters SubTrack Step-Size 0.01 3.0 5.0 15.0 10.0
Subspace Update Interval 500
Rank Config 8
$\alpha$ 2
BAdam Parameters Learning Rate 2E-05 2E-05 1E-05 2E-05 3E-05
Block Switch Interval 100
Switch Mode Random
BoolQ CB COPA WIC WSC AXg
Shared Parameters Batch Size 16
# Epochs 30
Learning Rate 1e-5
Max Seq. Len. 512
Low-Rank Optimizer SubTrack++ Step-Size 0.1 1.0 50.0 100.0 1.0 1.0
Methods Parameters Subspace Update Interval 500 100 100 500 250 100
Rank Config. 8
$\alpha$ 4
BAdam Parameters Block Switch Interval 100 50 50 100 50 50
Switch Mode Random
Appendix F Memory and Time Comparison
Table 8 presents the the peak memory consumption measured to compare SubTrack++ and other baselines. It shows that SubTrack++ requires on-par or better memory compared to GaLore [Zhao et al., 2024] and Fira [Chen et al., 2025] while getting state-of-the-art results. As detailed in Table 2, all baselines except BAdam [Luo et al., 2024] use the same number of optimizer parameters; therefore, any differences in their peak memory consumption stem from variations in their runtime parameters.
Table 8: Peak memory consumption comparison when pre-training Llama-based architectures with different sizes on C4 dataset. The 7B models are trained using the 8-bit Adam optimizer, except for the runs marked with . SubTrack++ demonstrates better or on-par memory compared to other low-rank methods that allows full-parameter training.
60M 130M 350M 1B 3B 7B
r=128 r=256 r=256 r=512 r=512 r=1024
Full-Rank 16.86 25.32 28.67 18.83 34.92 50.50
BAdam [Luo et al., 2024] 13.34 20.01 16.45 9.18 14.75 22.35
GaLore [Zhao et al., 2024] 16.89 25.52 27.85 14.74 26.03 36.00
LDAdam [Robert et al., 2025] 17.18 25.94 28.03 18.45 32.57 OOM
Fira [Chen et al., 2025] 16.39 24.99 27.33 14.68 25.62 47.84
SubTrack++ (Ours) 16.40 25.06 27.42 15.22 25.54 49.82
Table 9: Wall-time comparison for pre-training LLaMA-based architectures of varying sizes on the C4 dataset. The number of iterations is set to ensure 10 subspace updates for methods using periodic subspace updates: 2k for models ranging from 60M to 3B, and 5k for the 7B model. SubTrack++ achieves the lowest wall-time among all baselines that support full-parameter training on large models. The 7B models are trained using the 8-bit Adam optimizer, except for the runs marked with . Since this can impact wall-time comparisons, it is more appropriate to compare runs within the same cluster.
60M 130M 350M 1B 3B 7B
r=128 r=256 r=256 r=512 r=512 r=1024
Full-Rank 524.0 1035.1 1396.4 974.9 1055.9 12726.2
BAdam [Luo et al., 2024] 511.3 779.2 961.6 798.6 1004.1 7283.7
GaLore [Zhao et al., 2024] 547.8 1094.2 1589.0 1729.5 2715.5 21590.4
LDAdam [Robert et al., 2025] 639.9 1342.2 2083.4 2780.9 4625.4 OOM
Fira [Chen et al., 2025] 635.3 1180.8 1729.7 1938.5 2898.4 22554.3
SubTrack++ (Ours) 627.6 1140.6 1593.2 1304.3 1517.6 16491.7
Appendix G Pre-Training LLama-Based Architectures
We pre-trained all six Llama-based architectures for 10,000 iterations using hyperparameters reported in Table 10. Hyperparameters are selected based on reported values in GaLore paper [Zhao et al., 2024].
@@ Appendix G Pre-Training LLama-Based Architectures
SubTrack++ Step-Size 10

### v3

@@ 4 Experiments and Results
GaLore (Zhao et al., 2024) and Fira (Chen et al., 2025) periodically perform SVD to estimate the underlying subspace, while LDAdam (Robert et al., 2025) relies on the faster PowerSGD to update the subspace at every iteration. In contrast, SubTrack++ employs Grassmannian-based subspace tracking at the same frequency as GaLore and Fira. Comparing the time complexities of these methods, highlights why SubTrack++ is significantly more efficient than SVD-based methods. A breakdown of the subspace update time complexity for SubTrack++ is shown in Appendix D. Additionally, the memory required for storing optimizer states in SubTrack++ , is equivalent to GaLore and other baselines.
(a) Loss vs. Subspace Update Frequency.
(b) Loss vs. Subspace Update Rank.
Figure 4: Ablation results on (a) update frequency: decreasing the update interval (i.e., increasing the frequency) improves evaluation performance up to a point, but overly frequent updates hinder training convergence. (b) update rank: increasing the rank of updates degrades model performance, and beyond a certain threshold, can prevent convergence. These results emphasize the importance of controlled subspace adjustments.
Robust Subspace Tracking. Relying on SVD for subspace updates makes methods sensitive to noise and abrupt changes (He et al., 2025). Figure 6 compares Grassmannian subspace tracking with GaLore’s SVD on the Ackley function, highlighting how SVD causes erratic jumps, while our subspace tracking ensures robust optimization. GaLore struggles to reach the global minimum within 100 steps at scale factor 1, and although increasing the scale factor to 3 improves performance, it amplifies jumps that hinder convergence in non-convex settings, revealing sensitivity to hyperparameters, noise, and abrupt changes. To empirically measure the robustness of SubTrack++ in tracking evolving subspaces, we quantified subspace drift using the norm of the tangent vector $\nabla F$, which reflects the deviation between the projected gradient and the original gradient matrix. On Llama-350M pre-training, this norm rapidly decays from 0.06 to below 0.0002 within 5-7 subspace updates and remains near zero thereafter, indicating highly stable and well-aligned subspace tracking. This quantitative stability supports our qualitative observations in Figure 6 and aligns with prior findings in Grassmannian optimization.
5 Related Works
Parameter-Efficient Training. Several works aim to improve the efficiency of training LLMs, addressing a growing demand as their popularity rapidly increases. Popular LoRA (Hu et al., 2021) significantly reduces memory requirements for fine-tuning LLMs by leveraging two low-rank trainable low-rank matrices. Dettmers et al. (2024) employ quantization techniques and paged optimizers to further reduce memory usage. Additionally, Yaras et al. (2024) introduce Deep LoRA to address overfitting issues and reducing the need for precise tuning of the rank parameter. Several other works have also extended LoRA to enhance the efficiency of training and fine-tuning LLMs (Lialin et al., 2023; Renduchintala et al., 2024; Xia et al., 2024; Pan et al., 2024). Miles et al. (2024) propose compressing intermediate activations and reconstructing them during backpropagation to enhance memory efficiency. Yen et al. (2025) propose adjustments to the LoRA factorization that promotes balanced training and correspondingly update the optimizer’s internal states (i.e., first and second moments) to remain consistent under the change of basis. Additionally, Hao et al. (2024) demonstrate that full-parameter fine-tuning is feasible by random projections on the gradient matrix, showing that LoRA essentially performs a down-projection of the gradient. BAdam (Luo et al., 2024) leverages the block coordinate descent framework to reduce memory consumption while maintaining capabilities comparable to Adam.
@@ 5 Related Works
Gradient Low-Rank Projection. Several approaches aim to reduce optimizer states, as optimizers like Adam (Kingma and Ba, 2017) account for a significant portion of memory footprint (Li et al., 2023; Anil et al., 2019; Lv et al., 2024; Dettmers et al., 2022). MicroAdam (Modoranu et al., 2024) tackles this by compressing the gradient space and utilizing the compression error through feedback loops. Adam-mini (Zhang et al., 2024) partitions model into blocks, assigning a single learning rate to each block to preserve performance while saving memory. Gur-Ari et al. (2018); Schneider et al. (2024); Yaras et al. (2023) suggest that a substantial portion of gradients lies within a largely consistent subspace. GaLore (Zhao et al., 2024) leverages this fact to reduce the optimizer’s memory by projecting gradients into a low-rank subspace and then projecting them back for full parameter tuning. This approach has been integrated with other methods regarding efficient LLM training (Li et al., 2024). However, not all layers’ gradients evolve within a stable low-rank subspace. Jaiswal et al. (2024) identify layers where gradients evolve within a low-dimensional subspace and fine-tune only those layers, freezing the others to avoid inefficient low-rank updates. Grass (Muhamed et al., 2024) reduces memory usage by applying sparse projection matrices to the gradient. Ramesh et al. (2024) dynamically select and update a subset of parameters, for a fast and memory-efficient training. Fira (Chen et al., 2025) utilize a norm-based scaling method along with GaLore to maintain performance comparable to full-rank training. GoLore (He et al., 2025) addresses GaLore’s convergence issues and employ random projection in latter steps as a solutions. LDAdam (Robert et al., 2025) performs optimization within lower-dimensional subspaces, incorporating a projection-aware optimization update rule and a generalized error feedback mechanism. Projection-Aware APOLLO (Zhu et al., 2025) approximates channel-wise learning rate scaling based on random projection. Also, Liang et al. (2024) introduce a dynamically evolving projection matrix updated via online PCA, enhancing the model’s ability to navigate the parameter space efficiently without relying on expensive SVD.
@@
Geometric Subspace Updates. A common approach in working with high-dimensional data is to project the data into a lower-dimensional space, and many studies focus on tracking these subspaces as they evolve. Balzano et al. (2011) introduce an incremental method for updating subspaces on the Grassmannian when the data is partially observed. Zhang and Balzano (2016) and Kasai (2017) propose methods to handle noise effect in tracking these subspaces. Furthermore, Blocker et al. (2023) present a method for evolving geodesic-based data in the Grassmannian for updating the subspace effectively. Mo et al. (2025) propose LORO, a low-rank pretraining method that performs Riemannian optimization on the manifold of fixed-rank matrices, enabling parameter- and memory-efficient training by updating low-rank factors via manifold-aware gradients and retractions.
@@ 6 Discussion and Conclusion
Acknowledgments and Disclosure of Funding
Sirisha Rambhatla would like to acknowledge support of the Natural Sciences and Engineering Research Council of Canada (NSERC) Discovery Grant, RGPIN-2022-03512.
@@ Appendix A Convergence of SubTrack++
By applying vecvec to the gradient form given in the theorem, we obtain:
@@ Appendix C Projection-Aware Optimizer
Suppose the subspace changes at step $t$, transitioning from the subspace spanned by the orthonormal matrix $S_{t-1}$ to that spanned by $S_{t}$. Since both matrices are orthonormal—preserved by the Grassmannian update rule in (5)—the matrix $S_{t}^{\top}S_{t-1}$ represents the change of basis between the two subspaces. In other words, if $\mathbb{E}_{t-1}=(e_{t-1}^{1},\dots,e_{t-1}^{r})$ and $\mathbb{E}_{t}=(e_{t}^{1},\dots,e_{t}^{r})$ are orthonormal bases for the subspaces at steps t1t-1 and $t$, respectively, then the $i$th column of matrix $X$ transforms under the change of basis as $X^{i}_{t}=\sum_{j=1}^{r}\langle e^{i}_{t},e^{j}_{t-1}\rangle X^{j}_{t-1}$, where $X^{j}_{t-1}$ is the $j$th column of $X$ based on the basis of the time step t1t-1.
@@ Appendix D Time Complexity Analysis
Table 3 presents the time complexity breakdown for the subspace update step in the SubTrack++ algorithm assuming a $m\times n$ gradient matrix and rank $r$ projection matrix, where $r\ll m\leq n$. As outlined in Algorithm 1, the subspace update step begins by solving the least squares problem (2) to estimate the optimal update for $S_{t}$, the $m\times r$ orthonormal matrix. This operation has a time complexity of $O(mr^{2})$. Computing the residual and the partial derivative with respect to $S_{t}$ requires $O(mrn)$ and $O(mnr)$ time respectively. This is because the solution to the least squares problem, $A$, has shape $r\times n$ which is multiplied by $S_{t}$ in the residual $R=G_{t}-S_{t}A$, resulting in time complexity $O(mrn)$. The following operation for the partial derivative is $2RA^{T}$, where the matrix multiplication has $O(mnr)$ complexity. The tangent vector computation (4) which involves an identity transformation and matrix multiplication has time complexity of $O(m^{2}r)$.
@@
Rank-11 approximation of $\Delta F$ $O(mr^{2})$
@@
Appendix E Pre-Training Llama-Based Architectures
We pre-trained all six Llama-based architectures for 10k iterations using hyperparameters reported in Table 4. To demonstrate the generalizability of the proposed method, we also present results from pre-training the 7B architecture with both SubTrack++ and GaLore for 100k iterations, as shown in Figure LABEL:fig:7b-long. SubTrack++ maintains its advantage in terms of faster convergence and superior performance. The hyperparameters of this run are identical to those reported in Table 4, except for the number of iterations which is 100k. Also the bar-chart version of Figure 1 is represented in Figure 7.
@@ Appendix E Pre-Training Llama-Based Architectures
SubTrack++ Step-Size 10000